# Parallel computations for Metropolis Markov chains with Picard maps

### v1

@@ 5 Approximate Picard
By Theorem 1 the probability of correctly guessing each individual increment for the Picard map remains small up to $i=\mathcal{O}(d)$ steps. However, when accumulating all these small probabilities across $i$, the first incorrect guess happens much earlier, at $i=\mathcal{O}(\sqrt{d})$, see Corollary 1. This creates a bottleneck and does not allow us to leverage effectively more than $K=\mathcal{O}(\sqrt{d})$ parallel processors, see Theorem 2.
This motivated us to propose an Approximate Online Picard algorithm, which can leverage effectively up to $K=\mathcal{O}(d)$ processors, but whose output does not match anymore the output of the recursion in (1) and for which the limiting distribution of the resulting Markov chain has a small bias relative to $\pi$.
5.1 Approximate Online Picard algorithm
@@
is the proportion of mistakes made by the $j$-th Picard recursion in $\{L_{r}^{(j)}+1,\dots,\ell\}$, see also the pseudocode in Appendix C. Here $r$ acts as a tolerance parameter which controls the approximation error, and the case $r=0$ coincides with the (exact) Online Picard algorithm. By construction $L_{r}^{(j)}$ is non-decreasing w.r.t. $r$, and thus increasing $r$ reduces the runtime. Interestingly, for any $r>0$, the Approximate Online Picard algorithm is able to exploit up to $K=\mathcal{O}(d)$ parallel processors effectively, as shown in the following Proposition.
Proposition 4.
Under Assumptions 2-3, for all $\epsilon>0$, $r\in(0,1]$, $d\geq\log(5/(\epsilon r))$, $j\geq 1.5\log(3/(\epsilon r))$ and $K\leq\epsilon rd/(3c_{0})$, we have
@@
Proposition 4 shows that the fraction of incorrect guesses can be controlled with high probability up to $K=\mathcal{O}(d)$ processors and with only $j=\mathcal{O}(1)$ parallel iterations. This suggests that the Approximate Online Picard algorithm with $r>0$ and $K=\mathcal{O}(d)$ reduces the parallel run-time of the Online Picard algorithm by a factor $\mathcal{O}(\sqrt{d})$, thus converging in $j=\mathcal{O}(1)$ parallel iterations, as numerically verified in Section 6.
However, for $r>0$ the output no longer matches exactly the output of the sequential algorithm. Intuitively, we expect the error introduced to be small for small values of $r$, since the increments are coherent with the ones of the original RwM at least $(1-r)$-fraction of the times, so that, the Markov chain at those steps follows the correct transition kernel in (1). Also, the size of the other increments is bounded by the size of $Z$, which has Gaussian tails. The error of the Approximate Picard algorithm on the invariant distribution is numerically assessed in Section 6 for tolerance levels $r$ ranging from 0% to 20%, while a rigorous analysis of the error introduced by taking $r>0$ is outside the scope of this work.
@@ 6 Numerical simulations
We assess numerically the performance of the (Approximate) Online Picard algorithm applied to RWM and MwG Markov chains for high-dimensional regression models and the SIR epidemic model. The former simulations allow to test performances systematically and assess the tightness of the bounds developed in Section 3. The latter experiment illustrates the breadth of our results in a more sophisticated Bayesian problem where the posterior density is not log-concave and is only piecewise continuous. In this setting, state-of-the-art gradient-based methods such as Langevin and Hamiltonian Monte Carlo (Roberts and Tweedie, 1996a; Neal et al., 2011) cannot be directly used and zeroth-order (gradient-free) methods become a natural choice.
6.1 Algorithmic specification and performance metrics
In all numerical simulations, we set $\xi=c^{\mathrm{(RWM)}}/\sqrt{d}$ for RWM, and innovation $\tilde{Z}=\mathcal{N}(0,c^{\mathrm{(MwG)}})$ for MwG (with standard basis). The stepsizes $c^{\mathrm{(RWM)}}$ and $c^{\mathrm{(MwG)}}$ are tuned to obtain an average Metropolis acceptance-rejection of approximately 23.4% and 40%, respectively, as prescribed by the literature (Gelman et al., 1997, 1996). The code to reproduce the simulations is available at https://github.com/SebaGraz/ParallelMH.git.
@@ 6.1 Algorithmic specification and performance metrics
$\hat{G}=L^{(J)}/J\,,$
where $J$ is the total number of iterations of the (Approximate) Online Picard algorithm. We can interpret $\hat{G}$ as the multiplicative speedup obtained by parallel computing when ignoring overheads due to parallel execution.
@@
6.2 High-dimensional regressions
We consider three high-dimensional regression models:
  • E1
[…]
  • E2
[…]
  • E3
[…]
Figure 3 and Figure 4 summarize the results. The left panels display $\hat{G}$ as a function of $d$ (on a log-log scale), when $K=d$. The growth matches the one predicted by the theory of Section 3, i.e. $\mathcal{O}(\sqrt{d})$ for Online Picard and $\mathcal{O}(d)$ for the approximate version. The middle panels display $\hat{G}$ as a function of $K$ for fixed $d$ ($d=500$ for linear regression, $d=200$ otherwise). Again, the results align with the theory, with a linear speedup up to $K=\sqrt{d}$ (blue vertical dashed line) followed by a slower increase until $K=d$ and no increase afterwards. The right panel displays $\mathcal{M}_{r}$ and $\mathcal{E}_{r}$, defined in (12), as a function of $r$, for $K=d$ and $d=200$. Interestingly, for RWM, increasing $r$ tends to inflate the variance of the resulting approximate samples, in a way analogous to the bias observed in unadjusted Langevin algorithms (Roberts and Tweedie, 1996b; Dalalyan and Riou-Durand, 2020), while this is not the case for MwG, which appears to be significantly more robust to larger values of $r$. For these experiments, we set $N=10^{5}$ for the linear regression and $N=4\times 10^{5}$ for the logistic and Poisson regressions.
@@ 6.3 Convergence in the tails
We numerically assess the convergence of the Picard map for RWM chains started in the tails of the target distribution. Proposition 2 shows instantaneous convergence for log-concave targets in the limit for $|X_{0}|\to\infty$. Figure 5 numerically shows the convergence of $L^{(1)}$ to $K=200$ for initializations $X_{0}\sim\text{Unif}(x\in\mathbb{R}^{d}\colon\|x-x^{\star}\|=s)$, as $s$ gets larger, where $x^{\star}$ is the posterior mode and the target is given by the logistic model E2 in Section 6.2 with $d=200$ and a standard Gaussian prior. In all experiments, the tuning parameter $h$ was kept fixed and tuned in a preliminary run with the algorithm in stationarity. The results are averaged over 10 simulations.
Figure 3: Performance of Online Picard algorithm ($\bar{X}$) and its approximate versions ($\bar{X}_{r},\,r=0\%,\dots,20\%$) applied to RWM and MwG ($\bar{X}_{MwG}$), with target being the linear regression model E1. Left panel: average speedup $\hat{G}$ ($y$-axis, on a log-scale) with $K=d$, $N=10^{4}$ and $d=10^{2},\dots,10^{3}$ ($x$-axis, on a log-scale). Dashed lines $d\mapsto\sqrt{d}$ (blue) and $d\mapsto 3d/20$ (red) for reference. Middle panel: average speedup $\hat{G}$ ($y$-axis, on a log-scale) for $d=500$, $N=10^{4}$ and $K=2,\dots,10^{3}$ ($x$-axis, on a log scale). Vertical dash lines for $K=\sqrt{d}$ and $K=d$ for reference. Right panel: $\mathcal{M}$ and $\mathcal{E}$ as in (12) for the Approximate Picard algorithms with tolerance ($x$-axis) ranging from $0\%$ (exact Picard algorithm) to $20\%$ for $d=10^{2},\,K=d,\,N=10^{5}$.
Figure 4: Same as Figure 3 for the logistic regression model E2 (top panels) and the Poisson regression model E3 (bottom panels), with $N=10^{4},K=d$ (left panels), $N=10^{4},d=200$ (middle panels), $d=200,K=d,N=4\times 10^{5}$ (right panels).
Figure 5: $L^{(1)}$ ($y$-axis) for $X_{0}\sim\text{Unif}(x\in\mathbb{R}^{d}\colon\|x-x^{\star}\|=s)$, $s$ ranging from 0 to 2000 ($x$-axis) and target given by logistic regression E2 with $d=K=200$.
@@ 6.4 SIR epidemic model
$y_{i}(t)$ =$=\begin{cases}S&\text{if $i$ is \emph{susceptible} at time $t$}\\ I&\text{if $i$ is \emph{infected} at time $t$}\\ R&\text{if $i$ is \emph{removed} at time $t$}\\ \end{cases},$ $i=1,2,\dots,M.$
@@
Let $d\leq M$ be the total number of individuals that have been infected during the epidemic. We assume that the epidemic has originated with 1 initial infected individual and has been observed until the epidemic is extincted, that is, after the last infected infected individual has been removed.
[…]
@@
with $T=\max(x^{\circ}_{1},x^{\circ}_{2},\dots,x^{\circ}_{d})$ and $\mathcal{S}_{t}=\sum_{i=1}^{M}\mathbf{1}(y_{i}(t)=S)$, is the total number of person-to-person units of infectious pressure exerted during the course of the epidemic, see (Neal and Roberts, 2005, Section 5) for details of the likelihood and posterior measure in the set-up considered here.
@@
We simulate synthetic data and adopt the same experimental setting of Neal and Roberts (2005, Section 5.2), that is we set $M=200$, $\beta=0.001$ and $\gamma=0.15$. With this setting, the reproduction number is equal to $R=4/3$. The forward model generates $d=98$ infected individuals before the epidemic was extincted.
We set weak uninformative priors for $\beta$ and $\gamma$ with parameters $\lambda_{\beta}=\lambda_{\beta}=0.001$ and $\nu_{\beta}=\nu_{\beta}=1$ and run the Online-Picard applied to both RMW and MwG with target density as in (14) and for $N=10^{6}$ steps and with $K=\lfloor\sqrt{d}\rfloor=9$ parallel processors. The chain is initialized out of stationarity, at $X_{0,i}\sim x^{\circ}_{i}-\text{Exp}(0.05)$ (verifying that the posterior evaluated at this initial value is non-zero, that is $\pi(X_{0})\neq 0$). Moreover, we perform a second and computationally more intensive numerical experiment with $M=400,\,R=8/3,\,d=372,\,N=5\times 10^{6},\,K=\lfloor\sqrt{d}\rfloor=19$.
Table 1 shows the results obtained, where $\mathcal{M}_{0},\mathcal{E}_{0}$ were computed as in (12) relative to the $d$ dimensional latent observation $x$ with Markov chains with burn-in equal to $N/2$, while the true means and variances where estimated with a long run ($N=10^{7}$ for the first numerical experiment and $N=5\times 10^{7}$ for the second one, burn-in equal to $N/5$ in both cases) of RWM, started at the true values $x$. Figure 6 shows the functionals $B(X,x^{\circ})$ and $A(X,x^{\circ})$ as in (15), evaluated at every 10 steps of the RWM chain, with time index displayed with a yellow-blue gradient-color. The right panel, shows the corresponding values of $(\beta,\gamma)$, simulated from the conditional posteriors in (15).
Table 1: Online Picard applied to SIR model with $d=98(K=9,\,N=10^{6}$) and $d=372\,(K=19,\,N=5\times 10^{6})$.
Markov Chain $\hat{G}$ $\mathcal{M}_{0}$ $\mathcal{E}_{0}$
RWM $d=98$ 4.31 0.24 0.20
MwG $d=98$ 4.64 0.15 0.17
RWM $d=372$ 5.34 0.27 0.21
MwG $d=372$ 9.58 0.12 0.09
Figure 6: $(A(X_{i},x^{\circ}),B(X_{i},x^{\circ}))$ (left panel) and independent samples of $(\gamma,\beta)\mid(X_{10\,i},x^{\circ})$ (right panel), from $i=0$ (yellow) to $i=10^{5}$ (blue) and $X_{i}$ being the $i$th step of the RWM. The regions $R\geq 1$ is displayed in sky-blue. The true values from which we simulated the data are shown with red points.

### v2

@@ 5 Approximate Picard
By Theorem 3.1 the probability of correctly guessing each individual increment for the Picard map remains small up to $i=\mathcal{O}(d)$ steps. However, when accumulating all these small probabilities across $i$, the first incorrect guess happens much earlier, at $i=\mathcal{O}(\sqrt{d})$, see Corollary 3.4. This creates a bottleneck and does not allow us to leverage effectively more than $K=\mathcal{O}(\sqrt{d})$ parallel processors, see Theorem 3.5.
This motivates us to propose an Approximate Online Picard algorithm, which can leverage effectively up to $K=\mathcal{O}(d)$ processors, but whose output does not match anymore the output of the recursion in (1) and for which the limiting distribution of the resulting Markov chain has a small bias relative to $\pi$.
@@
is the proportion of mistakes made by the $j$-th Picard recursion in $\{L_{r}^{(j)}+1,\dots,\ell\}$, see also the pseudocode in Appendix D. Here $r$ acts as a tolerance parameter which controls the approximation error, and the case $r=0$ coincides with the (exact) Online Picard algorithm. By construction, $L_{r}^{(j)}$ is non-decreasing in $r$, and thus increasing $r$ reduces the runtime. Interestingly, for any $r>0$, the Approximate Online Picard algorithm is able to exploit up to $K=\mathcal{O}(d)$ parallel processors effectively, as shown in the following Proposition.
Proposition 5.1.
Under Assumptions 3.1-3.2, for all $\epsilon>0$, $r\in(0,1]$, $d\geq\log(5/(\epsilon r))$, $j\geq 1.5\log(3/(\epsilon r))$ and $K\leq\epsilon rd/(3c_{0})$, we have
@@
Proposition 5.1 shows that the fraction of incorrect guesses can be controlled with high probability up to $K=\mathcal{O}(d)$ processors and with only $j=\mathcal{O}(1)$ parallel iterations. This suggests that the Approximate Online Picard algorithm with $r>0$ and $K=\mathcal{O}(d)$ reduces the parallel run-time of the Online Picard algorithm by a factor $\mathcal{O}(\sqrt{d})$, thus converging in $j=\mathcal{O}(1)$ parallel iterations, as numerically verified in Section 6.
However, for $r>0$ the output no longer matches exactly the output of the sequential algorithm. Intuitively, we expect the error introduced to be small for small values of $r$, since the increments are coherent with the ones of the original RWM at least (1r)(1-r)-fraction of the times, so that the Markov chain at those steps follows the correct transition kernel in (1). Also, the size of the other increments is bounded by the size of $Z$, which has Gaussian tails. The error of the Approximate Picard algorithm on the invariant distribution is numerically assessed in Section 6 for tolerance levels $r$ ranging from 0% to 20%, while a rigorous analysis of the error introduced by taking $r>0$ is outside the scope of this work.
@@ 6 Numerical simulations
6.1 Overview
We assess numerically the performance of the (Approximate) Online Picard algorithm for high-dimensional regression models and the SIR epidemic model. The first simulations allow to test performance systematically and assess the tightness of the bounds developed in Section 3. The second experiment illustrates the breadth of our results for a more sophisticated Bayesian problem where the posterior density is not log-concave and is only piecewise continuous. In this setting, state-of-the-art gradient-based methods such as Langevin and Hamiltonian Monte Carlo (Roberts & Tweedie, 1996a; Neal et al., 2011) cannot be directly used and zeroth-order (gradient-free) methods become a natural choice. In this example, we also examine Discontinuous Hamiltonian Monte Carlo (D-HMC) (Nishimura et al., 2020), a non-reversible zeroth-order variant of MwG. Although our complexity bounds do not apply to D-HMC, the Online Picard algorithm applies directly because its Picard map is piecewise constant. In Section 6.8 we implement a parallel version of the online Picard algorithm on a concrete real-word application in precision medicine where zeroth-order parallel methods are appropriate (see Section G.1 for pseudocode and implementation details). For this application, the log-target can only be evaluated through a black-box computational routine that is expensive and does not provide gradients. Practical aspects such as hardware and software considerations, as well as parallelization overheads are discussed in detail in Section G in the supplement. See also Glatt-Holtz et al. (2024); Lee et al. (2010) for related discussions.
6.2 Algorithmic specification
In all numerical simulations, we set $\xi=c^{\mathrm{(RWM)}}/\sqrt{d}$ for RWM, and innovation $\tilde{Z}=\mathcal{N}(0,c^{\mathrm{(MwG)}})$ for MwG (with standard basis). The stepsizes $c^{\mathrm{(RWM)}}$ and $c^{\mathrm{(MwG)}}$ are tuned to obtain an average Metropolis acceptance-rejection of approximately 23.4% and 40%, respectively, as prescribed by the literature (Gelman et al., 1997, 1996). We set the number of integration steps for D-HMC to $L=3$ and the total number of iterations to $\lfloor N/(dL)\rfloor$, where $N$ is the total number of steps in RWM and MwG, ensuring that the total number of density evaluations is comparable across algorithms. To guarantee ergodicity, the stepsize of D-HMC is drawn as $h=|z|$ with $z\sim\mathcal{N}(0,c^{\mathrm{(MwG)}})$ (see Nishimura et al., 2020, Section 4.2). The code to reproduce the simulations is available at https://github.com/SebaGraz/ParallelMH.git.
6.3 Performance metric
@@
$\hat{G}=L^{(J)}/J\,$, (9)
where $J$ is the total number of parallel iterations of the (Approximate) Online Picard algorithm. $\hat{G}$ can be interpreted as the multiplicative speedup obtained by parallel computing when ignoring overheads due to parallel execution. In particular, we expect $\hat{G}$ to match the theoretical results in Theorems 3.5 and 4.1 and Proposition 5.1, that is $\hat{G}=\mathcal{O}(\sqrt{d})$ for the Online Picard algorithm with $K=\mathcal{O}(\sqrt{d})$ and $\hat{G}=\mathcal{O}(d)$ for the Approximate Picard algorithm with $K=\mathcal{O}(d)$. Moreover, we verify numerically that increasing $K$ beyond $\mathcal{O}(d)$ yields no further gains, as suggested by Theorem 3.1. The gains obtained by increasing $K$ from $\mathcal{O}(\sqrt{d})$ to $\mathcal{O}(d)$ for the Online Picard algorithm are more case-dependent, see details in Section 6.4 below.
Note that $\hat{G}$ is only a measure of speedup relative to the original sequential MCMC method. To evaluate the overall performance of a parallel MCMC method, the value of $\hat{G}$ should be combined with (i) a measure of statistical efficiency of the original MCMC, such as the Effective Sample Size (ESS), and (ii) the wall-clock time required to run the algorithm. In particular, when target evaluations are inexpensive (as in the toy regressions considered in Section 6.4), the overhead induced by parallelization (e.g., latency from synchronization and communication) may dominate, slowing the Picard algorithm. In Section 6.7 we report ESS values for the algorithms considered, and in Section 6.8 we report wall-clock times for parallel implementations of the Picard algorithm which also account for parallelization overheads.
6.4 High-dimensional regressions
We consider three high-dimensional regression models:
  • E1
    Linear regression: for $i=1,2,\dots,n$, $Y_{i}\mid A_{i}\overset{\text{i.i.d.}}{\sim}\mathcal{N}(\langle A_{i},x\rangle,\sigma^{2}),\sigma^{2}=1$ and unknown parameter $x\in\mathbb{R}^{d}$. The sample size is set to $n=5d$.
  • E2
    Logistic regression: for $i=1,2,\dots,n$, $Y_{i}\mid A_{i}\overset{\text{i.i.d.}}{\sim}\mbox{Ber}(\psi(\langle A_{i},x\rangle)),$ with $\psi(u)=(1+e^{-u})^{-1}$ and unknown parameter $x\in\mathbb{R}^{d}$. The sample size is set to $n=10d$.
  • E3
    Poisson regression: for $i=1,2,\dots,n$, $Y_{i}\mid A_{i}\overset{\text{i.i.d.}}{\sim}\mbox{Poiss}(\exp(\langle A_{i},x\rangle)),$ with unknown parameter $x\in\mathbb{R}^{d}$. The sample size is set to $n=10d$.
In all cases, we set independent standard Gaussian priors for the parameters. With this setting, the corresponding posteriors are analytical for the linear regression, while not tractable for the logistic and Poisson regression. Furthermore, for the latter, the gradient of $\log\pi$ decays exponentially fast in the tails, making standard first-order Monte Carlo methods numerically unstable. We generate covariates as $A_{i}\overset{\text{i.i.d.}}{\sim}\mathcal{N}_{d}(0,I_{d}/d)$, some true parameters $x\sim\mathcal{N}_{d}(0,I_{d})$ and $Y_{1:n}$ from the corresponding model. We run multiple times our algorithms, each time varying the number of processors $K$ and the parameter dimension $d$. We initialize the algorithms at the posterior mean for linear regression and at the true parameter value otherwise. Unless specified otherwise, we run algorithms until $L^{(J)}\geq N=10^{4}$. Figure 3 and Figure 4 summarize the results for RWM.
The left panels display $\hat{G}$ as a function of $d$ (on a log-log scale), when $K=d$ and $K=\sqrt{d}$. The growth rates match the ones predicted by the theory of Section 3, i.e. $\mathcal{O}(\sqrt{d})$ for Online Picard and $\mathcal{O}(d)$ for the approximate version. Note that increasing $K$ from $\sqrt{d}$ to $d$ in the Online Picard algorithm improves $\hat{G}$ only by a constant factor, indicating that, while there might be a practical benefit in setting $K$ beyond $\sqrt{d}$, the rate in Theorem 3.5 is tight for RWM.
The right panels display $\hat{G}$ as a function of $K$ for fixed $d$ ($d=500$ for linear regression, $d=200$ otherwise). Again, the results align with the theory, with a linear speedup up to $K=\sqrt{d}$ (blue vertical dashed line) followed by a slower increase until $K=d$ and no increase afterwards. Similar experiments are performed for MwG and are available in the supplementary material, Appendix E. For MwG, we generally notice better parallelization performance compared to RWM. In particular, choosing $K$ larger than $\sqrt{d}$ appears to yield more than a constant improvement in speedup. This may be due to the nearly isotropic targets considered in our regression problems. In such settings, the Online Picard algorithms for MwG can achieve optimal speedup as suggested by Proposition 4.2.
Figure 3: Performance of Online Picard algorithm ($\bar{X}$) and its approximate versions ($\bar{X}_{r},\,r=5\%,\dots,20\%$) applied to RWM, with target being the linear regression model E1. Left panel: average speedup $\hat{G}$ ($y$-axis, on a log-scale) with $K\in\{\sqrt{d},d\}$, $N=10^{4}$ and $d=10^{2},\dots,10^{3}$ ($x$-axis, on a log-scale). Dashed lines $d\mapsto\sqrt{d}$ (blue) and $d\mapsto 3d/20$ (red) for reference. Right panel: average speedup $\hat{G}$ ($y$-axis, on a log-scale) for $d=500$, $N=10^{4}$ and $K=2,\dots,1500$ ($x$-axis, on a log scale). Vertical dashed lines for $K=\sqrt{d}$ and $K=d$ for reference.
Figure 4: Same as Figure 3 for the logistic regression model E2 (top panels) and the Poisson regression model E3 (bottom panels), with $N=10^{4},K\in\{\sqrt{d},d\}$ (left panels). $N=10^{4},d=200$ (right panels).
6.5 Bias of Approximate Online Picard
@@
Figure 5 displays $\mathcal{M}_{r}$ and $\mathcal{E}_{r}$, defined in (10), as a function of $r$, for $K=d$ and $d=100$. When $r=0$, the Approximate Online Picard algorithm reduces to the standard Online Picard algorithm, and $\mathcal{M}_{0},\mathcal{E}_{0}$ correspond to the Monte Carlo estimation error arising from taking a finite number of steps $N$. Interestingly, for RWM, increasing $r$ tends to inflate the variance of the resulting approximate samples, in a way analogous to the bias observed in unadjusted Langevin algorithms (Roberts & Tweedie, 1996b; Dalalyan & Riou-Durand, 2020), while this is not the case for MwG, which appears to be significantly more robust to larger values of $r$. For these experiments, we set $N=10^{5}$.
Figure 5: $\mathcal{M}$ and $\mathcal{E}$ as in (10) for the Approximate Picard algorithms with tolerance ($x$-axis) ranging from $0\%$ (exact Picard algorithm) to $20\%$ for $d=10^{2},\,K=d,\,N=10^{5}$. $N=4\times 10^{5}$ (right panels) for the linear regression E1 (left panel), logistic regression E2 (middle panel), Poisson regression E3 (right panel). Dashed lines for MwG.
@@ 6.6 Convergence in the tails
We numerically assess the convergence of the Picard map for RWM and MwG chains started in the tails of the target distribution. Figure 6 numerically shows the value of $L^{(1)}$, averaged over 10 simulations, for $K=200$, $d=200$, random initializations $X_{0}\sim\text{Unif}(x\in\mathbb{R}^{d}\colon\|x-x^{\star}\|=s)$, with $x^{\star}$ being the posterior mode, and target distribution given by posterior distribution of the logistic model E2 in Section 6.4. In all experiments, the tuning parameter $h$ was kept fixed and tuned in a preliminary run with the algorithm in stationarity. Our results suggest that $L^{(1)}$ converges to $K$ as $s$ increases, which is coherent with Proposition 3.7.
Figure 6: Empirical averages of $L^{(1)}$ ($y$-axis) for $X_{0}\sim\text{Unif}(x\in\mathbb{R}^{d}\colon\|x-x^{\star}\|=s)$, $s$ ranging from 0 to 2000 ($x$-axis) and target given by logistic regression E2 with $d=K=200$.
@@ 6.7 SIR epidemic model
[…]
@@
Let $d\leq M$ be the total number of individuals that have been infected during the epidemic. We assume that the epidemic has originated with 1 initial infected individual and has been observed until the epidemic becomes extinct, that is, after the last infected individual has been removed.
As it is common for this class of models, we do not observe the infection times (i.e. when the coordinate process changes from $S\to I$) and we only observe the times where infected individuals change their state from $I$ to $R$. Hereafter, we denote by $x\in\mathbb{R}^{d}$ the latent vector of infection times and by $x^{\circ}\in\mathbb{R}^{d}$ with $x_{i}\leq x^{\circ}_{i},\,1\leq i\leq d$, the observed removed times of the infected population. We assume that $\gamma_{0},\beta_{0}$ are unknown. The likelihood for $(\gamma_{0},\beta_{0})$ becomes tractable only when the infection times of each infected individual are known and is given by
@@
with $T=\max(x^{\circ}_{1},x^{\circ}_{2},\dots,x^{\circ}_{d})$ and $\mathcal{S}_{t}=\sum_{i=1}^{M}\mathbf{1}(y_{i}(t)=S)$, is the total number of person-to-person units of infectious pressure exerted during the course of the epidemic, see (Neal & Roberts, 2005, Section 5) for details of the likelihood and posterior measure in the set-up considered here.
@@
$\gamma\mid x\sim\Gamma(d+\nu_{\gamma},\,1/(\lambda_{\gamma}+B(x,x^{\circ}))),\qquad\beta\mid x\sim\Gamma(d+\nu_{\beta}-1,\,1/(\lambda_{\beta}+A(x,x^{\circ})))$ (13)
We simulate synthetic data and adopt the same experimental setting of Neal & Roberts (2005, Section 5.2), that is we set $M=200$, $\beta=0.001$ and $\gamma=0.15$. With this setting, the reproduction number is equal to $R_{0}=4/3$. The forward model generates $d=98$ infected individuals before the epidemic went extinct.
We set weak uninformative priors for $\beta$ and $\gamma$ with parameters $\lambda_{\beta}=\lambda_{\beta}=0.001$ and $\nu_{\beta}=\nu_{\beta}=1$ and run the Online-Picard applied to RWM, MwG and D-HMC (Nishimura et al., 2020) with target density as in (12) and for $N=10^{6}$ steps and with $K=\lfloor\sqrt{d}\rfloor=9$ parallel processors. The chain is initialized out of stationarity, at $X_{0,i}\sim x^{\circ}_{i}-\text{Exp}(0.05)$ (verifying that the posterior evaluated at this initial value is non-zero, that is $\pi(X_{0})\neq 0$). Moreover, we perform a second and computationally more intensive numerical experiment with $M=400,\,R_{0}=8/3,\,d=372,\,N=5\times 10^{6},\,K=\lfloor\sqrt{d}\rfloor=19$.
Table 1 shows the results obtained. Here $\hat{G}$ is computed as in (9), $\mathcal{M}_{0}$ and $\mathcal{E}_{0}$ as in (10), and $d$ denotes the dimensionality of $x$. The Effective Sample size, which is defined in Section F of the supplementary material, is estimated using the effectiveSize function of the coda R package (Plummer et al., 2006), applied to each coordinate of the Online Picard output (i.e. the full trajectory after burn-in). We report both the average and the minimum across coordinates. These values are complementary to those of $\hat{G}$ (as they do not directly relate to the parallelization performance of the Picard algorithm) and provide an empirical basis for comparing the three MCMC methods under consideration.
@@
Table 1: Online Picard applied to SIR model with $d=98$ ($K=9,\,N=10^{6}$) and $d=372$ ($K=19,\,N=5\times 10^{6}$). Under ESS, we report the average ESS and minimum ESS in brackets. Bold characters for the best performance.
[…]
RWM $d=98$ 4.28 0.22 0.25 36.6 (4.69) 156.65 (20.07)
MwG $d=98$ 4.94 0.09 0.12 404.57 (151.81) 1998.58 (749.94)
D-HMC $d=98$ 3.98 0.08 0.09 508.15 (207.60) 2022.44 (826.25)
RWM $d=372$ 5.15 0.39 0.26 18.58 (1.62) 44.18 (13.90)
MwG $d=372$ 9.70 0.12 0.10 169.68 (22.61) 1645.89 (219.32)
D-HMC $d=372$ 7.40 0.11 0.08 270.05 (42.13) 1998.37 (311.76)
Figure 7: $(A(X_{i},x^{\circ}),B(X_{i},x^{\circ}))$ (left panel) and independent samples of $(\gamma,\beta)\mid(X_{10\,i},x^{\circ})$ (right panel), from $i=0$ (yellow) to $i=10^{5}$ (blue) and $X_{i}$ being the $i$th step of the RWM. The regions $R_{0}\geq 1$ is displayed in sky-blue. The true values from which we simulated the data are shown with red points.
6.8 A real-world application for zeroth-order parallel algorithms
We apply here a parallel implementation of the online Picard algorithm on a real statistical application in precision medicine. In this application, each patient measurement is modeled as a complex delayed ordinary differential equation, and the goal is to infer the model parameters to understand the effect of personalized treatments on advanced-stage cancer patients, see Jenner et al. (2021) for more details. This application provides a concrete example where, due to black-box log-target evaluations where gradients are not available, zeroth-order MCMC methods are a natural choice. This example is also well-suited to parallel methods, given that each target evaluation is computationally expensive, as it requires running numerical solvers.
A full description of the model and dataset is provided in (Gagnon et al., 2023, Section 5.2), where it was also used as a test-case for parallel zeroth-order MCMC methods. The parameter dimension is $d=14$. On an Apple M3 CPU with 1616GB of RAM, a single evaluation of $\pi$ takes approximately $0.25$ seconds, while one iteration of a parallel implementation of the Picard algorithm with $K=8$ parallel CPU cores takes approximately $0.4$ seconds, see Section G in the supplement for details. We set the total number of iterations to $N=10^{5}$. For this application, we observe $\hat{G}=4.37$ and an effective speedup (i.e. reduction in total wall-clock time, see Section G.2) of 2.52 when using the parallel implementation of the online Picard algorithm rather than the sequential implementation of RWM. These results are consistent with (31) in Section G.2 and the theoretical results in Section 3.